OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 3x − 7 3x − y = −2

OpenStudy (anonymous):

The answer choices are (3, 11) and (−3, −7) (11, 3) and (−3, −7) (3, 11) and (−7, −3) No real solutions

OpenStudy (anonymous):

I would solve using substitution, since you are already given what y equals in the first equation. You know y = x^2 +3x -7, so you can plug that value of y into the second equation giving you: 3x - y = -2 3x - (x^2 + 3x -7) = -2 -x^2 + 3x - 3x +7 = -2 -x^2 +7 = -2 -x^2 = -9 x^2 = 9 x = (+-)sqrt(9)

OpenStudy (anonymous):

now that you've solved for x, you can plug that value back into either original equation to solve for y! Give it a shot!

OpenStudy (anonymous):

so it would be y=\[\sqrt{9}^{2}+3\sqrt{9}-7\]

OpenStudy (anonymous):

so its plus or minus square root of 9, so you can simplify that to -3 and +3. but otherwise yes, thats perfect!

OpenStudy (anonymous):

Thanks